### v3

# MLEM: Generative and Contrastive Learning as Distinct Modalities for Event Sequences

Abstract
This study explores the application of self-supervised learning techniques for event sequences. It is a key modality in various applications such as banking, e-commerce, and healthcare. However, there is limited research on self-supervised learning for event sequences, and methods from other domains like images, texts, and speech may not easily transfer. To determine the most suitable approach, we conduct a detailed comparative analysis of previously identified best-performing methods. We find that neither the contrastive nor generative method is superior. Our assessment includes classifying event sequences, predicting the next event, and evaluating embedding quality. These results further highlight the potential benefits of combining both methods. Given the lack of research on hybrid models in this domain, we initially adapt the baseline model from another domain. However, upon observing its underperformance, we develop a novel method called the Multimodal-Learning Event Model (MLEM). MLEM treats contrastive learning and generative modeling as distinct yet complementary modalities, aligning their embeddings. The results of our study demonstrate that combining contrastive and generative approaches into one procedure with MLEM achieves superior performance across multiple metrics.
**footnotetext: These authors contributed equally to this work
1 Introduction
Motivation. Event Sequences (EvS) are widely used in various real-world applications, including medicine Waring et al. (2020), biology Valeri et al. (2023), social medial analysis Liguori et al. (2023), fault diagnosis  Shao et al. (2019), churn prediction  Jain et al. (2021); Sudharsan and Ganesh (2022), customer segmentation  Carnein and Trautmann (2019), fraud detection  Xie et al. (2022) and more Zhuzhel et al. (2023); Fursov et al. (2021). As a result, there is a growing need to effectively model such data. Most self-supervised methods are focused on images, text, speech, or time-series. In Computer Vision (CV), contrastive learning methods have achieved state-of-the-art results as a pre-training strategy before supervised fine-tuning He et al. (2021); Chen et al. (2020); Caron et al. (2021). In Natural Language Processing (NLP), most modern models rely on generative pre-training Radford and Narasimhan (2018); Devlin et al. (2019).
Self-Supervised Learning (approaches have not been extensively explored for EvS. However, applying existing SSL methods directly to EvS may be unreasonable or even impossible due to several key differences between EvS and other well-studied domains: a. Multiple Features, events in EvS can be described by multiple features, which adds complexity to the learning process. b. Categorical and Numerical Values, event features in EvS may include both categorical and numerical values, further complicating modeling. c. Non-Uniform Temporal Spacing, events in EvS are not uniformly spaced in time, which requires models to handle irregular temporal patterns. d. Temporal Proximity, it is to assumed that events occurring in close temporal proximity share relevant information. Given these unique characteristics of EvS, further research is needed to explore SSL approaches specifically tailored to ES and address these challenges. We describe a formal structure of EvS in Section 2.
The aim of our work is to answer the following questions: Q1: Which self-supervised approach is superior for ES modeling: generative or contrastive? Q2: Can we efficiently combine generative and contrastive approaches into one procedure to obtain more robust results?
The Contributions. To answer these questions, we (1) study established generative, contrastive, and hybrid approaches. (2) To the best of our knowledge, we are the first to examine generative pre-training, for EvS classification, and hybrid methods in the domain of EvS in general. (3) We introduce a new technique named MLEM, inspired by contemporary multimodal studies, such as (Radford et al., 2021; Zhai et al., 2023; Moskvoretskii and Kuznetsov, 2023). MLEM allows for the combination of various SSL approaches by treating differently pre-trained models as distinct modalities. More precisely, MLEM aligns the contrastive and generative embeddings. We demonstrate that such combination achieves superior performance on average across multiple datasets and metrics compared to existing methods.
Evaluation employs a diverse set of datasets to cover various aspects of the problem space: three real-world transactional datasets from the SeqNAS benchmark (Udovichenko et al., 2024), the widely-used medical dataset Physionet 2012, and a simulated dataset of pendulum motion designed to examine explicit time dependencies. This selection ensures that our findings are robust across different types of data and tasks. To comprehensively measure performance we employ a set of metrics, including downstream tasks of regression or classification of sequences, prediction of the next event type or timing, intrinsic dimension, and anisotropy.
Baselines. To ensure a clear and fair comparison, we have selected only the top-performing model from each setting for our analysis, comparison of baseline approaches is presented in Table 3 and Table 3. These include models from recent studies on contrastive learning (Babaev et al., 2022), generative modeling in EvS (Lin et al., 2022), and hybrid approaches from other domains Kim et al. (2021). We focus on these models because they are clearly superior in their categories, making the inclusion of lesser models unnecessary.
Figure 1: An example of one sequence $S_{i}$ with some categorical and real-valued sub-events $\{k^{1},k^{2},\ldots\}$.
We provide the source code for all the experiments described in this paper.***The source code of our work is publicly available at https://github.com/VityaVitalich/MLEM.
2 Event Sequences Preliminaries
Datasets used in this work can be described as sets of sequences: $C=\{(S_{1},y_{1}),(S_{2},y_{2}),\ldots\}$, where $y_{i}$ is an attribute corresponding to the entire sequence or some target. Each $S_{i}=((t^{1}_{i},\Upsilon^{1}_{i}),(t^{2}_{i},\Upsilon^{2}_{i}),\ldots)$ is a sequence of events $x^{j}_{i}=(t^{j}_{i},\Upsilon^{j}_{i})$, where $t_{i}^{j}$ represents the time when the event $x_{i}^{j}$ occurred and $t_{i}^{j}\leq t^{j+1}_{i}$. The set of sub-events $\Upsilon^{j}_{i}=\{k^{1},k^{2},\ldots\}$ describes the event $x_{i}^{j}$. It is important to note that in this work, the terms Event Sequences (EvS) and Irregularly Sampled Time-series (ISTS) are used interchangeably, as the occurrence of measurement or sampling for the latter can be seen as an event. Figure 1 depicts an example of one sequence.
Figure 2: Schematic Models Overview: A) The Contrastive Model divides sequence $S_{i}$ into subsequences $\{S_{i}^{\prime},S_{i}^{\prime\prime}\}$, encodes them into latent vectors $\{h_{i}^{\prime},h_{i}^{\prime\prime}\}$, and applies contrastive loss $L_{\text{con}}$. B) The Generative Model autoregressively reconstructs sequence $S$ cross-attended to latent vector $h$, derived from $S$ via a bottleneck encoder. Language model loss $L_{\text{LM}}$ is applied to reconstructed sequence $\hat{S}$. C) The MLEM Hybrid Model integrates generative and contrastive approaches, treating them as distinct modalities. It generates latent vectors $h_{i}^{g}$ and $h_{i}^{c}$ from a learned generative encoder and frozen, pretrained contrastive encoder respectively. An alignment loss $L_{\text{align}}$ is utilized to align vectors from the same sequence while separating those from different sequences. Language modeling loss $L_{\text{LM}}$ is applied as in generative model.
3 Methodology
3.1 Contrastive Learning
For our contrastive learning approach, we employ COLES Babaev et al. (2022) framework, as it has shown superior performance over other contrastive methods. It learns to encode the sequence $S$ into a latent vector $h\in\mathcal{R}^{m}$ by bringing sub-sequences of the same sequence closer in the embedding space, while pushing sub-sequences from different sequences further apart. The overall procedure is illustrated in Figure 2 and formal definition is given in Appendix Contrastive learning
3.2 Generative Modeling
For our generative model, we consider the best performing setting from GNTPP (Lin et al., 2022) and adapt it to our task.
For the encoder, we utilize RNN, as it shows dominance to our task in recent studies (Udovichenko et al., 2024). It extracts the final hidden state $h$, which is subsequently passed into a Transformer decoder (Vaswani et al., 2017). The decoder autoregressively reconstructs the entire sequence, conditioned on $h$ via the Cross-Attention mechanism. This entire process is illustrated in Figure 2.
To train the model, we use the next event prediction objective, which is similar to training language models. Nonetheless, as each event in the sequence $x^{j}_{i}=(t^{j}_{i},\Upsilon^{j}_{i})$ contain multiple sub-events $\Upsilon=\{k^{1},k^{2},\ldots\}$, we need to reconstruct all of them. To this end, we apply the cross-entropy loss for each categorical $k$ and the Mean Squared Error (MSE) loss for each real-valued $k$. The final loss is a cumulative sum of the losses for all elements in $\Upsilon=\{k^{1},k^{2},\ldots\}$, plus the MSE loss for time $t$. We denote this loss as $L^{LM}$.
Intuitively, this procedure requires our encoder to develop a representation informative enough for the decoder to accurately reconstruct the entire sequence. Since we map all our sequences $S$ into, $h\in\mathcal{R}^{m}$ we call all encoders bottleneck encoders.
Table 1: EvS modeling and choice of baselines
Supervised Contrastive Generative
Paper SeqNAS CoLES Generative TPP
2024 2022 2022
Best base models GRU GRU REV-ATT, LSTM
Table 2: Hybrid SSL approaches
Sum of Losses (Naïve) Multimodal
Paper DINO 2023,
iBOT 2022 MLEM
GCRL 2021
Domain Images and Text EvS
Table 3: Evaluation of methods fine-tuned with supervised loss for downstream task.
Method ABank Age PhysioNet Pendulum TaoBao
ROC-AUC Accuracy ROC-AUC MSE ROC-AUC
Supervised $0.768\pm 0.000$ $0.602\pm 0.005$ $0.790\pm 0.021$ $\underline{0.33\pm 0.02}$ $0.684\pm 0.002$
Contrastive $0.729\pm 0.033$ $0.638\pm 0.007$ $\mathbf{0.815\pm 0.013}$ $\mathbf{0.26\pm 0.02}$ $0.679\pm 0.003$
Generative $\underline{0.788\pm 0.003}$ $0.639\pm 0.007$ $0.787\pm 0.007$ $\mathbf{0.26\pm 0.03}$ $\mathbf{0.695\pm 0.004}$
Naïve $0.658\pm 0.020$ $0.638\pm 0.007$ $0.759\pm 0.024$ $\mathbf{0.26\pm 0.04}$ $0.691\pm 0.002$
MLEM $\mathbf{0.790\pm 0.000}$ $\mathbf{0.642\pm 0.005}$ $0.780\pm 0.004$ $\mathbf{0.26\pm 0.05}$ $\mathbf{0.695\pm 0.002}$
3.3 Naïve Hybrid Method
We consider weighted summation of contrastive and generative loss as baseline hybrid model, as it showed dominant performance in previous studies (Kim et al., 2021). We call this approach Naïve since it is a straightforward combination of losses $L=\alpha L^{LM}+\beta L^{con}$.
3.4 MLEM
In our method, named MLEM, we propose to combine contrastive learning and generative modeling by treating them as distinct modalities with further alignment. To align the embeddings obtained through generative modeling with those acquired through contrastive learning, we employ a contrastive aligning procedure inspired by CLIP Radford et al. (2021). We utilize the recent SigLIP loss Zhai et al. (2023) for this purpose.
Overall MLEM training procedure is nearly the same as for the generative model described in Section 3.2, except the alignment, which incurs an additional loss (1) resulting in the total loss (2). Unlike CLIP and other multimodal models, our approach involves pre-training only one modality. Specifically, we pre-train the contrastive encoder first and then train the generative encoder, aligning it with the fixed contrastive encoder.
Formally, both generative $\mathcal{E}_{g}$ and contrastive $\mathcal{E}_{c}$ encoders receive a set of sequences $C=\{S_{1},S_{2},\ldots\}$ and map each sequence $S$ into corresponding latent vectors $h^{g}$ and $h^{c}$. The goal of MLEM alignment is to pull embeddings obtained from the same sequence closer to each other and to push embeddings from different sequences further away. This procedure is illustrated in Figure 2. Below is the alignment loss we use:
$L^{align}=\frac{1}{|C|}\sum_{i=1}^{|C|}\sum_{j=1}^{|C|}\underbrace{\log\frac{1}{1+e^{z_{ij}(-t\cdot{h^{g}_{i}}\cdot h^{c}_{j}+b)}}}_{L^{align}_{ij}}$ (1)
Here, $z_{ij}\in\{-1,1\}$ indicates if a pair of embeddings originated from the same sequence $S$, $t$ and $b$ denote the temperature and bias, respectively, both are learnable parameters.
Total loss to train MLEM:
$L=\alpha L^{LM}(S,\hat{S})+\beta L^{align}(\mathcal{E}_{g}(S),\mathcal{E}_{c}(S))$ (2)
Here, $S$ and $\hat{S}$ denote original and reconstructed sequences, $\alpha$ and $\beta$ are hyperparameters that adjust the strength of alignment.
4 Experiments
4.1 Datasets
In our study, we use five established EvS datasets formally outlined in Section 2.
  • A subset of EvS datasets from SeqNAS benchmark Udovichenko et al. (2024), more precisely: ABank, Age, TaoBao. ABank and Age are bank transactions, and TaoBao is user activity.
  • PhysioNet 2012 comprises highly sparse medical event data, with the goal of predicting mortality Silva et al. (2012).
  • Pendulum is a synthetic dataset simulating pendulum motion with the objective of predicting its length using a sequence of coordinates. This dataset was created specifically to evaluate a model’s capability to effectively incorporate the time component in sequence modeling.
The detailed description is provided in Appendix Detailed Dataset Information.
Table 4: Evaluation of self-supervised methods using linear and non-linear probing for downstream and TPP tasks. The best-performing values are highlighted in bold, while the second-best values are underlined.
ABank Age PhysioNet Pendulum TaoBao
ROC-AUC TPP Acc. Acc. TPP Acc. ROC-AUC TPP MSE MSE TPP MSE ROC-AUC TPP Acc.
Linear Probing
Contrastive $0.678\pm 0.003$ $0.37$ $0.623\pm 0.010$ $0.28$ $\mathbf{0.819\pm 0.003}$ $1.05$ $\mathbf{0.37\pm 0.02}$ $0.8$ $0.680\pm 0.001$ $0.22$
Generative $0.753\pm 0.003$ $0.43$ $0.610\pm 0.004$ $0.3$ $0.759\pm 0.010$ $\mathbf{0.93}$ $0.74\pm 0.09$ $0.8$ $\mathbf{0.689\pm 0.005}$ $\mathbf{0.35}$
Naïve $0.703\pm 0.002$ $0.37$ $0.602\pm 0.003$ $\mathbf{0.31}$ $0.733\pm 0.005$ $\mathbf{0.93}$ $0.59\pm 0.07$ $0.8$ $0.680\pm 0.003$ $0.32$
MLEM $\mathbf{0.757\pm 0.000}$ $\mathbf{0.43}$ $\mathbf{0.627\pm 0.000}$ $0.29$ $0.762\pm 0.010$ $\mathbf{0.93}$ $0.41\pm 0.01$ $\mathbf{0.79}$ $0.683\pm 0.003$ $\mathbf{0.35}$
Non-linear Probing
Contrastive $0.691\pm 0.004$ $0.37$ $0.629\pm 0.010$ $0.28$ $\mathbf{0.822\pm 0.001}$ $1.05$ $\mathbf{0.37\pm 0.00}$ $0.8$ $0.686\pm 0.001$ $0.22$
Generative $0.758\pm 0.003$ $0.43$ $0.618\pm 0.004$ $0.3$ $0.764\pm 0.006$ $\mathbf{0.93}$ $0.63\pm 0.02$ $0.8$ $0.684\pm 0.002$ $\mathbf{0.35}$
Naïve $0.704\pm 0.005$ $0.37$ $0.608\pm 0.002$ $\mathbf{0.31}$ $0.774\pm 0.008$ $\mathbf{0.93}$ $0.57\pm 0.05$ $0.8$ $\mathbf{0.690\pm 0.002}$ $0.32$
MLEM $\mathbf{0.759\pm 0.003}$ $\mathbf{0.43}$ $\mathbf{0.634\pm 0.005}$ $0.29$ $0.780\pm 0.001$ $\mathbf{0.93}$ $0.40\pm 0.01$ $\mathbf{0.79}$ $0.688\pm 0.002$ $\mathbf{0.35}$
4.2 Evaluation objectives
To evaluate the effectiveness of self-supervised training strategies, we follow previous studies Dubois et al. (2024, 2021) and focus on two key aspects:
  • Fine-Tuning Performance. The downstream metrics after fine-tuning the entire network.
  • Embeddings Probing. The quality of the embeddings for downstream and TPP tasks tested via linear and non-linear probing.
As well, we include additional objectives for embedding quality, that were discovered in previous studies Razzhigaev et al. (2023); Nakada and Imaizumi (2020) , that are connected with performance:
  • Anisotropy. It assesses the non-uniformity of embeddings in space, providing insights into the contextualization of our embedding. Lower anisotropy in embeddings has been linked to better model performance Ait-Saada and Nadif (2023).
  • Intrinsic dimension. It evaluates the optimal dimensionality of data, showing information captured by the embeddings.
Objectives details are described in appendix Objectives description.
4.3 Models
Base model and Feature embeddings. To transform features into a numerical vector, we use an embedding layer for categorical features and linear projection for numerical ones. For ABank, Age and TaoBao we set the dimension of each feature to 32. For Pendulum and PhysioNet dimension is set to 8. Additionally, we use the time difference between events as a separate feature.
We use GRU as an encoder because it has been proven effective in encoding time-ordered sequences  Rubanova et al. (2019); Tonekaboni et al. (2021); Yoon et al. (2019); Udovichenko et al. (2024). Udovichenko et al. performs a Neural Architecture Search on various models and highlights that the presence of RNN blocks in the architecture is associated with higher performance on average. Lin et al. illustrate that the performance gain brought by history encoders is very small for TPP modeling. Based on these empirical observations, we believe that our design choice is reasonable.
We it with a single-layer and a hidden size of 512, we take the last hidden state as sequence embedding. For consistency in our experiments, we use the same hyperparameters across all training strategies and datasets.
For Supervised model, we use the aforementioned encoder with a linear head.
Contrastive learning approach is described at 3.1 and uses the same encoder as mentioned above. Furthermore, we integrate a Feed Forward Projector atop the GRU for loss calculation, a technique commonly adopted in several studies for enhanced performance Grill et al. (2020); Oquab et al. (2023).
For the generative modeling, we employ a vanilla transformer decoder configured with LayerNorm, consisting of 3 layers, 2 heads, a hidden size of 128, and a Feed Forward hidden size of 256. To ensure training stability, we also apply LayerNorm to the sequence embedding produced by the encoder. Additionally, a projection layer is used atop the decoder to predict each feature.
4.4 Training details
To maintain consistency in our experiment, we trained all models using identical parameters. Models were trained for 100 epochs on datasets with fewer than 100K sequences and for 40 epochs on larger datasets. The learning rate was set at 1e-3, weight decay at 3e-3, and the batch size was fixed at 128. For the Hybrid models, we set $\alpha$ at 1 and $\beta$ at 10. We averaged the results across three different seeds to ensure reliability and reproducibility. The standard deviation could be not reported if negligible.
5 Results
5.1 Fine-tuning performance
First, we evaluate the effectiveness of fine-tuning the entire model after pre-training procedure. The results presented in Table 3 indicate that, among all the evaluated pre-training methods, MLEM consistently achieves the most favorable results across all datasets, except for the highly sparse PhysioNet dataset. Furthermore, it is noteworthy that solely supervised learning consistently produces inferior results.
5.2 Embeddings probing
We conducted a comprehensive analysis of the quality of embeddings using both linear and non-linear probing techniques, results are presented in Table 4.
Our evaluation reveals that neither the contrastive nor the generative approach consistently outperforms the other. However, by effectively integrating the advantages of both contrastive and generative techniques, MLEM model demonstrates superior performance across most datasets on average. Notably, when MLEM is not the top performer, it is often ranked second, highlighting its versatility.
Table 5: Average anisotropy and intrinsic dimension across datasets
Method Anisotropy $\downarrow$ Intrinsic Dimension $\uparrow$
Contrastive $0.11\pm 0.04$ $10.15\pm 6.12$
Generative $0.08\pm 0.03$ $14.86\pm 10.12$
Naïve $0.07\pm 0.04$ $11.26\pm 6.51$
MLEM $\mathbf{0.06\pm 0.03}$ $\mathbf{15.86\pm 9.02}$
Table 6: Performance change in linear probing performance after shuffling events order, higher is better
Method ABank Age PhysioNet Pendulum TaoBao
Contrastive +0.47% -1.93% -0.37% -367.57% -0.15%
Generative -0.93% +0.16% -6.19% -91.89% -0.29%
Naive -1.42% +0.99% -3.00% -147.46% 0.00%
MLEM -1.45% -0.16% -6.56% -285.37% +0.29%
5.3 Anisotropy and Intrinsic Dimension
We conducted an evaluation of anisotropy and the intrinsic dimensions of embeddings from various pre-training strategies, as detailed in 4.2.
Results presented in Table 6 show that MLEM has the highest intrinsic dimension and the lowest anisotropy, on average, across all datasets. This finding may indicate that our pre-training strategy effectively enables the embeddings to encapsulate a greater amount of information and ensures their uniform distribution in space.
However, we found no correlation between intrinsic dimension or anisotropy and downstream performance in our domain, as illustrated in correlation plot 4(b). While our findings align with the most recent study by Dubois et al. (2024) regarding anisotropy, they differ in that they observed a correlation between intrinsic dimension and performance.
Figure 3: The t-SNE visualizations showcase embeddings resulting from various pre-training strategies. The left portion displays the embeddings for the Age dataset, while the bottom row illustrates those for the TaoBao dataset. Each point represents a sequence $S_{i}$ from a given dataset, colored accordingly to the corresponding attribute $y_{i}$. For Age, there are 4 classes and 2 classes for TaoBao
5.4 Visualization
The findings related to anisotropy and intrinsic dimension align with the t-SNE visualization of the Age and TaoBao datasets shown in Figure 3. While contrastive embeddings typically form distinct clusters, the generative approaches and MLEM provide a more complex embedding space.
6 Embedding Robustness
Our study also explores the robustness of model embeddings to specific types of perturbations. We focus on examining the impact of dropout and the shuffling of events within one sequence on the linear probing performance.
We assess robustness with classification or regression tasks. In each scenario, we assess either the class of the whole sequence or the real value describing the whole sequence (pendulum length). We refer to this assessment as global and assume that this attribute does not change over the entire sequence (e.g., user age group). Attributes that change over time are referred to as local (e.g., amount of payment or next purchase category).
6.1 Sensitivity to permutation
We randomly permuted order of events along time axis and analyzed performance of pre-trained models. Models were trained without permutation. The results, presented in Table 6, show that the sequence order has only a small effect on most of the results. This suggests that current models may be employing a ’Bag-of-Words’ approach, treating the events as a set rather than considering the sequential nature of the data when determining certain global attributes. Additionally, we observed that the dataset’s degradation is greater when it has fewer categorical features.
Despite the small degradation, it is still present, highlighting the need for modeling temporal relations. Furthermore, all the models demonstrate the ability to encode sequential data, as evidenced by a decline in performance on the pendulum dataset where the time component is crucial.
The assessment of the effect of order permutation on the ability to model local attributes is similar to the TPP task, and we leave it for the future work.
6.2 Sensitivity to data omission
To investigate the effects of data omission, we applied a random dropout strategy, removing entire events from a sequence along the time axis, with varying probabilities $p_{dropout}\in\{0.1,0.3,0.5,0.7\}$. We report the average percentage decrease in linear probing performance, depicted in Figure 4(a).
We notice that MLEM performance slightly decreases at $0.1$ dropout probability and more so at higher levels. This indicates that MLEM is more sensitive to dropout and requires the presence of all the samples rather than depending on a couple specific samples. This property can be used in a sensitivity analysis or related applications.
Additionally, the generative modeling embeddings exhibited an intriguing behavior: they not only resisted the decline, but actually surpassed their original performance at a dropout rate of 0.5. We believe that such effect show high sensitivity to the presence of noisy or out-of-distribution events. Nevertheless, to draw reliable conclusions, further work is required.
(a) Mean percentage change averaged across datasets. The X-axis represents the dropout probability, the Y-axis indicates the mean change in the linear probing, compared to the metric with no dropout. Error bars reflect 1 sigma deviation.
(b) Correlation plots for anisotropy and intrinsic dimension with normalized performance.
7 Computational complexity
To obtain the MLEM encoder, we need to train two models, which is not desirable. However, to demonstrate the practicality of our method, we compare MLEM with SeqNas Udovichenko et al. (2024) in terms of performance and computational efficiency.
Table 7: Comparison of downstream metrics and computational costs for training MLEM versus SeqNas.
Method ABank Age
ROC-AUC GPU hours Accuracy GPU hours
MLEM $0.790\pm 0.000$ $47$ $0.642\pm 0.005$ $9$
SeqNas $0.7963\pm 0.001$ $288$ $0.645\pm 0.002$ $80$
SeqNas employs an architectural search to identify the most optimal supervised architecture for the task at hand. We posit that SeqNas represents an upper bound for downstream metrics. Table 9 presents results from the two largest datasets, examined by us, demonstrating that MLEM achieves performance near the upper bound set by SeqNas, but with significantly lower computational cost. Detailed compute budget reported in table 9
8 Related work
In the following section, we will review various generative, contrastive, hybrid, and other related methods and works.
Generative methods for pre-training and representation learning have seen notable development in recent years in both NLP Radford and Narasimhan (2018); Devlin et al. (2019); Raffel et al. (2023) and CV He et al. (2021); Assran et al. (2023); Bachmann et al. (2022). In the EvS domain, the sole study by Lin et al. (2022) investigates various generative models for TPP tasks. Based on their findings, we have considered their best performing setting to serve as a generative model for our study, with slight modifications to fit our tasks.
Contrastive methods are a widely recognized in He et al. (2020); Grill et al. (2020). These methods typically draw closer the embeddings of variations of the same object and distance them from those of different objects. In CoLES Babaev et al. (2022) authors study contrastive learning for EvS. In their work, the positive samples consist of subsequences derived from a single user’s events, while the negative samples are obtained from events of different users. We utilize this method as best performing along contrastive approaches.
Hybrid methods. Our work falls into the category of hybrid self-supervised approaches Qi et al. (2023); Yang et al. (2023); Lazarow et al. (2017) Few studies have merged generative and contrastive methods, typically focusing on loss combination. DINOv2 Oquab et al. (2023) updates the traditional contrastive loss from DINO Caron et al. (2021) with iBOT’s Zhou et al. (2022) reconstruction loss through Masked Image Modeling. Similarly, GCRL Kim et al. (2021) combines these losses via a weighted sum, significantly outperforming models that are purely generative or contrastive. We follow the same methodology to form the hybrid baseline method with simply summing contrastive and generative losses.
Supervised learning is commonly used for the EvS classification task. The most comprehensive recent paper SeqNAS Udovichenko et al. (2024) adopts Neural Architecture Search to identify optimal architectures without strong prior assumptions about the underlying process. A key finding is the effectiveness of Recurrent Neural Networks (RNNs) in modeling EvS, leading us to incorporate RNN encoder in our model. Furthermore, the authors propose benchmark datasets for EvS classification. We selected the largest datasets from the benchmark, as self-supervised methods require substantial data volumes.
9 Limitations
Physionet underperformane. We observe consistent underperformance of methods that include generative approach on the PhysioNet dataset. This could be attributed to the high rate of missing values, which complicates the generation of accurate predictions for subsequent steps. Furthermore, the relatively small size of the dataset may be a contributing factor, given that generative modeling is generally dependent on large amounts of data.
Compute Resources. One limitation of our model is the requirement for a higher computational budget compared to previous methods. However, in the domain of event sequences, the primary challenge is typically achieving the highest metric rather than optimizing for computational efficiency.
Datasets and Models. We selected the most relevant datasets to cover a broad spectrum of tasks and chose the best-performing models from previous studies to represent SSL methods. Nonetheless, the comparison could be improved by incorporating a wider variety of datasets and model architectures.
Scalability. Although we aimed to ensure a fair comparison, the results may vary with different scales of data or models. However, observed trend show that more data leads to better results on MLEM and generative method.
10 Conclusion
To our knowledge, our study is the first to apply generative modeling to event sequences for representation learning and pre-training. We found that neither generative nor contrastive pre-training definitively outperforms the other.
In response, we developed the MLEM method, which creatively combines contrastive learning with generative modeling. Inspired by recent advances in multimodal research, MLEM treats these approaches as separate modalities and consistently outperforms existing models on most datasets across both primary and secondary objectives. This finding suggests a promising new direction for enhancing self-supervised learning across various domains.
Additionally, our research provides several key insights. Firstly, the best current self-supervised models do not utilize the sequential nature of data effectively. We also discovered that random event deletion does not adversely affect performance; in some cases, it actually enhances it. Secondly, unlike findings in other domains (Dubois et al., 2024; Ait-Saada and Nadif, 2023), we observed no correlation between anisotropy or intrinsic dimension and downstream performance.